# date21082024

### v1

@@ Abstract
Figure 2 depicts the variation of $\log[\det(\mathbf{S})]$ in the complex $\omega$ plane for the unperturbed network, i.e., $\Delta n=0$. Bright regions, marked with white dots and crosses, correspond to the poles of the unperturbed network scattering matrix. Figures 2(a)-(d) show detailed plots of the smaller regions bounded by the white dashed boxes in the main panel. The positions of a subset of the poles, specifically those marked with white crosses, were tracked as $\Delta n$ was increased. The trajectories followed are shown by the solid red and white lines emanating from the initial pole positions (crosses). White lines were traced by numerically solving $\det(\mathbf{S}^{-1})=0$ for each value of $\Delta n$, using the previous pole position as an initial guess. The red lines, on the other hand, were traced using Eq. (3) with $\alpha=\Delta n$ to determine the pole shifts at each step (see Supplemental Material [34] for details of the calculation of $\mathbf{Q}_{\alpha}$). As can be seen, the Wigner-Smith theory agrees excellently with the direct numerical solutions. Some discrepancies, however, between the two methods can be seen, e.g., near the top of the main panel. In this region, it was found that the computed values for $\mathop{\mathrm{Res}}_{\omega=\omega_{p}}\mbox{tr}(\mathbf{Q}_{\alpha})$ were relatively large, implying that the pole shifts were more sensitive to the perturbation. Achieving greater accuracy would thus require small a perturbation step size to maintain validity of Eq. (2). Dashed red and white lines emanating from the crosses on the right hand side of the figure show the pole trajectories when $\Delta n$ instead took negative values, and show further agreement between the two methods.
Figure 2 shows several interesting features. First, note that the half space $\mathrm{Im}(\omega)>0$ contains dark regions for which $\det(\mathbf{S})=0$. These zeros can be pushed to the real axis by introducing loss to the system, whereby they correspond to coherent perfect absorption modes [46]. Note also that the poles in our data exhibit two distinct types of trajectories: some follow long, meandering paths, while others revolve around localized, closed loops. Figures 2(a)-(b) in particular demonstrate the latter type. It is important to realize that $\mathbf{S}$ is periodic in $\Delta n$ since for specific values of $\Delta n$ the additional propagation phase acquired in $\Omega_{\alpha}$ will be an exact multiple of $2\pi$. At these values $\mathbf{S}$ will be identical to that of the unperturbed network, such that poles following loop trajectories must return to their original positions, while poles following open trajectories must pass through the positions of other poles of the unperturbed network (white dots in Figure 2).
@@
As a further simple application of our theory, we next considered adding gain to (or ‘pumping’) our network which can lead to random lasing. Mathematically, gain can be introduced to the $i$’th link by varying its refractive index $n_{i}$ according to $n_{i}=(n_{i})-i\gamma_{i}$, where $(n_{i})$ is calculated as before and $\gamma_{i}>0$ is a variable gain coefficient taken, for simplicity, to be frequency independent over the pumping bandwidth. Adding gain tends to cause poles to drift towards the real axis and the pole that first reaches the axis dictates the dominant lasing frequency [47]. Single mode lasing at arbitrary resonant frequencies is achievable by shaping the pump profile in accordance with the spatial profile of the resonant mode [48, 49]. Similar selective lasing can be realized using our theory by considering the collection of generalized Wigner-Smith operators associated with the gain coefficients. Specifically, for any pole $\omega_{p}$ and for all $i$ indexing the links, we can calculate $\mathop{\mathrm{Res}}_{\omega=\omega_{p}}\mbox{tr}(\mathbf{Q}_{\gamma_{i}})$, where $\mathbf{Q}_{\gamma_{i}}$ is the operator associated with adding gain $\gamma_{i}$ to only the $i$’th link. In light of Eq. (3), these values predict how the pole will shift when different links are pumped in isolation, thus revealing which link should be pumped to optimally shift $\omega_{p}$ towards the real axis.
@@
Figure 3 shows the results of the described numerical pumping experiment. The imaginary parts of three arbitrarily selected poles (distinguished by color and corresponding to the numbered poles in Figure 2) were tracked in response to three alternate pumping methods (distinguished by line style). Solid lines track the poles when all links were pumped uniformly, in which case the narrowest resonance, mode 1 (at 3424.47 THz), reaches the lasing threshold first. Alternatively, the dashed and dot-dashed lines track the poles when the links highlighted (and denoted $2$ and $3$ respectively) in Figure 1, were pumped in isolation. In particular, links 2 and 3 correspond to those predicted to optimally shift modes 2 (at 3424.8 THz) and 3 (at 3426.25 THz) respectively towards the real axis. This predicted behaviour is evident in Figure 3, with pumping of link 2 bringing mode 2 to lasing threshold first (and similarly for mode/link 3). We also calculated the spatial intensity profiles of modes 2 and 3 across the network, which are shown on the right hand side of Figure 3. Notably, the intensity of mode 2 (3) is strongly peaked across link 2 (3), confirming the wisdom that the optimal pump profile should conform to the mode’s spatial distribution. Interestingly, however, the use of the generalized Wigner-Smith operator eliminated the need to explicitly calculate the mode distribution in determining the pump profile. Finally, we note in passing that a similar strategy of selectively introducing loss to specific links could be used to achieve selective coherent perfect absorption of desired modes.
@@
Conclusion.—To conclude, in this work we have presented a novel method for calculating resonance shifts in perturbed open systems using generalized Wigner-Smith operators associated with the perturbation parameters. These operators have found increasing use in the control and manipulation of optical fields in recent years. Our work reveals the connection of generalized Wigner-Smith operators to resonant properties in non-Hermitian systems and further highlights their utility. Our perturbation theory is based on generic complex analytic arguments and is therefore applicable to a wide range of scenarios. At the same time, we have demonstrated that our results reduce to more traditional perturbation formulas for high quality resonances. We have verified our theory numerically by tracking pole shifts caused by refractive index perturbations in a complex photonic network, and in a spatially selective pumping experiment. Our work provides a novel way to analyze scattering resonances, which may be of use in cavity or nanostructure design [50] and future sensing technologies.
@@ II Derivation of the volume integral form of the Wigner-Smith operator
In this section we present a derivation of Eq. (5) in the main text, which expresses the Wigner-Smith operators in terms of volume integrals over the system.
@@
which is Eq. (5) in the main text.

### v2

@@ Abstract
Figure 2 depicts the variation of $\log[|\det(\mathbf{S})|]$ in the complex $\omega$ plane for the unperturbed network, i.e., $\Delta n=0$. Bright regions, marked with white dots and crosses, correspond to the poles of the unperturbed network scattering matrix. Figures 2(a)-(d) show detailed plots of the smaller regions bounded by the white dashed boxes in the main panel. The positions of a subset of the poles, specifically those marked with white crosses, were tracked as $\Delta n$ was increased. The trajectories followed are shown by the solid red and white lines emanating from the initial pole positions (crosses). White lines were traced by numerically solving $\det(\mathbf{S}^{-1})=0$ for each value of $\Delta n$. The red lines, on the other hand, were traced using Eq. (3) with $\alpha=\Delta n$ to determine the pole shifts at each step (see Supplemental Material [39] for details of the calculation of $\mathbf{Q}_{\alpha}$). As can be seen, the WS theory agrees excellently with the direct numerical solutions. Some discrepancies, however, can be seen, e.g., near the top of the main panel. In this region, it was found that computed values for $\mathop{\mathrm{Res}}_{\omega=\omega_{p}}\mbox{tr}(\mathbf{Q}_{\alpha})$ were relatively large, implying that the pole shifts were more sensitive to the perturbation. Accordingly, a smaller perturbation step size is requried to restore validity of Eq. (2). Dashed red and white lines emanating from the crosses on the right hand side of the figure show the pole trajectories for $\Delta n<0$ and show further agreement between the two methods.
Figure 2 shows several interesting features. First, note that the half space $\mathrm{Im}(\omega)>0$ contains dark regions for which $\det(\mathbf{S})=0$. These zeros can be pushed to the real axis by introducing loss to the system, whereby they correspond to coherent perfect absorption modes [52]. Note also that the poles in our data exhibit two distinct types of trajectories: some follow long, meandering paths, while others revolve around localized, closed loops. Figures 2(a)-(b) in particular demonstrate the latter type. It is important to realize that $\mathbf{S}$ is quasiperiodic in $\Delta n$ since for specific values of $\Delta n$ the additional propagation phase acquired in $\Omega_{\alpha}$ will be an exact multiple of $2\pi$. At these values, the only difference between the perturbed and unperturbed network will be the weak Fresnel matrices at the ends of the perturbed link segment, the effect of which is negligible for small $\Delta n$. Poles following loop trajectories must therefore return to their original positions, while poles following open trajectories must pass through the positions of other poles of the unperturbed network (white dots in Figure 2).
@@
As a further application of our theory, we next considered adding gain to (or ‘pumping’) our network which can lead to random lasing [47]. Mathematically, gain can be introduced to the $j$’th link by setting its refractive index to $n_{j}=(n_{j})-i\gamma_{j}$, where $(n_{j})$ is calculated as before and $\gamma_{j}>0$ is a gain coefficient taken, for simplicity, to be frequency independent over the pumping bandwidth. Adding gain causes poles to drift towards the real axis and the pole that reaches the axis first typically determines the dominant lasing frequency [53]. Single mode lasing at arbitrary resonant frequencies is achievable by shaping the pump profile in accordance with the spatial profile of the resonant mode [54, 55]. Similar selective lasing can be predicted using our theory by considering the collection of generalized WS operators associated with the gain coefficients. Specifically, for any pole $\omega_{p}$ and for $j$, we calculate $\mathop{\mathrm{Res}}_{\omega=\omega_{p}}\mbox{tr}(\mathbf{Q}_{\gamma_{j}})$, where $\mathbf{Q}_{\gamma_{j}}$ is the operator associated with adding gain $\gamma_{j}$ to only the $j$’th link. In light of Eq. (3), these values predict how the pole will shift when different links are pumped, thus revealing which should be pumped to optimally shift $\omega_{p}$ towards threshold.
@@
Figure 3 shows the results of the described numerical pumping experiment. The imaginary parts of three arbitrarily selected poles (distinguished by color and corresponding to the numbered poles in Figure 2) were tracked in response to three alternate pumping methods (distinguished by line style). Solid lines track the poles when all links were pumped uniformly, in which case the narrowest resonance, mode 1 (at 3424.47 THz), reaches the lasing threshold first. Alternatively, the dashed and dot-dashed lines track the poles when the links highlighted (and denoted $2$ and $3$ respectively) in Figure 1, were pumped in isolation. In particular, links 2 and 3 correspond to those predicted to optimally shift modes 2 (at 3424.8 THz) and 3 (at 3426.25 THz) respectively towards the real axis. This predicted behaviour is evident in Figure 3, with pumping of link 2 bringing mode 2 to lasing threshold first (and similarly for mode/link 3). We also calculated the spatial intensity profiles of modes 2 and 3 across the network, which are shown on the right hand side of Figure 3. Notably, the intensity of mode 2 (3) is strongly peaked across link 2 (3), confirming the wisdom that the optimal pump profile should conform to the mode’s spatial distribution. Interestingly, however, the use of the generalized WS operator eliminated the need to explicitly calculate the mode distribution in determining the pump profile. Finally, we note that a similar strategy of selectively introducing loss to specific links could produce coherent perfect absorption of desired modes.
@@
Conclusion—In this work we have presented a novel method for calculating resonance shifts in perturbed open systems using generalized WS operators. Our work reveals the connection between generalized WS operators and resonant properties in non-Hermitian systems and further highlights their utility. Our perturbation theory is based on general complex analytic arguments and is therefore applicable to a broad range of scenarios. Furthermore, we demonstrated our results reduce to traditional perturbation formulas for high quality resonances. We have verified our theory numerically by tracking pole shifts caused by refractive index perturbations in a complex photonic network, and in a spatially selective pumping experiment. Our work provides a novel way to analyze scattering resonances, which may be of use in cavity or nanostructure design [56] and future sensing technologies.
Acknowledgements.
N.B. was supported by Singapore Ministry of Education Academic Research Fund (Tier 1) Grant RG66/23. M.R.F. was supported by funding from the Institute for Digital Molecular Analytics and Science (IDMxS) under the Singapore Ministry of Education Research Centres of Excellence scheme and by Nanyang Technological University Grant SUG:022824-00001.
@@ II Derivation of the volume integral form of the Wigner-Smith operator
In this section we present a derivation of Eq. (Perturbing scattering resonances in non-Hermitian systems: a generalized Wigner-Smith operator formulation) in the main text, which expresses the Wigner-Smith operators in terms of volume integrals over the system.
@@
which is Eq. (Perturbing scattering resonances in non-Hermitian systems: a generalized Wigner-Smith operator formulation) in the main text.